The random sample shown below was selected from a normal distribution. 7,5,9,5,7,3 Complete parts a and b. a. Construct a 99% confidence interval for the population mean . (Round to two decimal places as needed.) on b. Assume that sample mean x and sample standard deviation s remain exactly the same as those you just calculated but that are based on a sample of n 25 observations. Repeat part a. What is the effect of increasing the sample size on the width of the confidence intervals? The confidence interval is). What is the effect of the sample size on the width of the confidence interval? 0 A. As the sample size increases, the width increases dcpe s n menvalsR Round to two decimal places as needed.) B. As the sample size increases, the width stays the same. O C. As the sample size increases, the width decreases.
Answer #1

Solution:-

a) 99% confidence interval for the population mean u is C.I = (2.54, 9.45).

C.I = \bar{x}\pm t_{\alpha /2}\ast \frac{\sigma }{\sqrt{n}}

C.I = 6 + 4.032*0.85635

C.I = 6 + 3.453

C.I = (2.547, 9.453)

b) (C) As the sample size increases, the width decreases.

99% confidence interval for the population mean u is C.I = ( 4.83, 7.17).

C.I = \bar{x}\pm t_{\alpha /2}\ast \frac{\sigma }{\sqrt{n}}

C.I = 6 + 2.797*0.41952

C.I = 6 + 1.1734

C.I = ( 4.8266, 7.1734)

As the sample size increases, the width decreases.